Grade 12Mechanics

a particle moves in xy plane under the action of force F such that the linear momentum at any time t is P(x)=2cos(t),P(y)=2sin(t) the angle theta b/w P & F @ a t is

guru
8 Years agoGrade 12

2 Answers

Karan Gupta
8 Years ago
P(x) = 2 cos(t), F(x) = dP(x)/dt = -2 sin(t)
P(y) = 2 sin (t), F(y) = dP(y)/dt = 2 cos (t)
 
Dot product of momentum and Force = F(x)*P(x) + F(y)*P(y) = -4 sin(t)cos(t) + 4 sin(t)cos(t) = 0
Since dot product is zero, angle between the two is 90 degrees
Kushagra Madhukar
5 Years ago
Dear student,
Please find the answer to your question.
 
P(x) = 2 cos(t), F(x) = dP(x)/dt = -2 sin(t)
P(y) = 2 sin (t), F(y) = dP(y)/dt = 2 cos (t)
 
Dot product of momentum and Force = F(x)*P(x) + F(y)*P(y) = -4 sin(t)cos(t) + 4 sin(t)cos(t) = 0
Since dot product is zero, angle between the two is 90 degrees
 
Thanks and regards,
Kushagra